One question, five parts to the same question.
Scribners Corporation produces fine papers in three production departments-Pulping. Drying and Finishing. In the Pulping DepaComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 Pre
Answer #1
1) Equivalent units of production
units % pulping % CC
Completed & transferred 152,000 100% 152,000 100% 152,000
work in process inventory,March 31 8,000 100% 8,000 75% 6000
Equivalent units of production 160,000 158,000
2) cost per Equivalent units
pulping CC total
Beginning work in process 1,584 1,024 2,608
cost added during the year 81,616 55,856 137,472
total cost 83,200 56,880 140,080
equivalent units 160,000 158,000
cost per EUP 0.52 0.36 0.88
3) Cost of ending work in proces
pulping 8,000 * 0.52 = 4160
Conversion 6000 * 0.36 = 2160
Cost of ending work in proces 6320 answer
4) Cost of units transferred
pulping 152,000 * 0.52 = 79040
conversion 152,000 * 0.36 = 54720
Cost of units transferred 133760 answer
5) CostReconcilation report
cost to be accounted for
cost of beginning work in process inventory 2,608
Cost added to production during the period 137,472
cost to be accounted 140,080
cost accounted for
cost of units completed & transferred out 133,760
Cost of ending work in process inventory 6320
Total cost accounted for 140,080